OpenStudy (anonymous):

Solve the Exponential equation Algebracally (16 + .878/26)3t=30

OpenStudy (anonymous):

Answer: t=0.007278458

OpenStudy (anonymous):

(16+(0.88)/(26))^(3)t=30 Divide each term in the equation by (16+(0.88)/(26))^(3). ((16+(0.88)/(26))^(3)t)/((16+(0.88)/(26))^(3))=(30)/((16+(0.88)/(26))^(3)) Cancel the common factor of (16+(0.88)/(26))^(3). t=(30)/((16+(0.88)/(26))^(3)) Simplify the right-hand side of the equation by simplifying each term. t=0.01

OpenStudy (anonymous):

hopefuly helps you out

OpenStudy (anonymous):

Thanks.